# Machine Learning Approaches to Vocal Register Classification in Contemporary Male Pop Music

### v1

@@ Abstract
For singers of all experience levels, one of the most daunting challenges in learning technical repertoire is navigating placement and vocal register in and around the passagio (passage between chest voice and head voice registers). Particularly in pop music, where a single artist may use a variety of timbre’s and textures to achieve a desired quality, it can be difficult to identify what vocal register within the vocal range a singer is using. This paper presents two methods for classifying vocal registers in an audio signal of male pop music through the analysis of textural features of mel-spectrogram images. Additionally, we will discuss the practical integration of these models for vocal analysis tools, and introduce a concurrently developed software called AVRA which stands for Automatic Vocal Register Analysis. Our proposed methods achieved consistent classification of vocal register through both Support Vector Machine (SVM) and Convolutional Neural Network (CNN) models, which supports the promise of more robust classification possibilities across more voice types and genres of singing.
@@ 1 Introduction
In all voices, there are two muscles responsible for manipulating the vocal folds. The thyroarytenoid (TA) muscles shorten and thicken the folds while the cricothyroid (CT) muscles lighten and lengthen. TA dominant vocal functions result in ”chest voice” which the lower speech like vocal register in all genders. On the other hand, the CT muscles produce that pop falsetto head voice sound that has become prominent in the 21st century, and is often used to reach higher notes with a softer quality.
Traditionally, ”mixed voice” is used to denote the coordination of both TA and CT muscles, which imbues the sound with qualities from both chest voice and head voice. However, since an experienced singer can change the balance of muscle function, the term mixed voice covers the entire spectrum of chest-like to head-like timbres. For this reason, we divided the mixed voice classification into ”mix” and ”head mix” where mix denotes a more TA dominant, solid sound and head mix reflects more of the lighter head voice qualities.
The challenge in identifying vocal register is the variety of factors involved from person to person. Vocal registers don’t have clear boundaries and can overlap in pitch particularly through the passagio, more commonly known as the vocal break. Besides pitch, qualities like resonance, and placement can change the color and texture of the sound. This is how a singer can sing one note in two very different ways which is reflected in their choice of vocal register.
Expressing all these qualities of vocal audio with data was done by rendering samples of a single vocal register into mel-spectrograms. Thus, the proposed methods were achieved by analyzing the features and properties of mel-spectrograms.
The related software application may help students and experienced singers further their understanding and practice of contemporary vocal technique. It could be viewed as a teaching aid since the processed output is applied to the spectrogram which may help singers pinpoint where their register shifts occur and visualize the qualities they are trying to replicate.
@@ 2.1 Audio Classification with Images
3 Methods
This section describes the systems used for building the dataset, pre-processing images and extracting features, and architecture of the models (primarily CNN). Finally, it will conclude by summarizing the training and evaluation process that determined which models were pushed to production.
@@ 3.1 Dataset
Similarly to the paper referenced earlier, we used mel-spectrograms as the expression of the labeled audio clips. The vocals from mp3 audio files of pop-songs were split using a tool called a ”stem splitter” in Logic Pro. Then the raw vocals were imported into Audacity and rendered into mel-spectrograms with a maximum frequency of 20000 Hz, gain of 20 dB, range of 80 dB, standard Roseus color scheme, and Hann window type as the default settings. These parameters were later replicated using the Librosa library in Python for the backend of the desktop application.
@@
Given that the dataset consists of 2010’s pop music, pitches ranged from C3-C5, but were generally from A3-G4. Qualitative definitions for each vocal register were agreed upon at the beginning of the project with supplementation from various vocal technique texts (Peckham 2010). Then the raw vocals of each song were split into 3 second clips and rendered into spectrograms. An image was captured for any continuously used vocal register and then later processed into a standard size for the model inputs. These sub-images were given a one-hot coded label for one of the 4 vocal registers. In these labels, 0 corresponds to Chest Voice, 1 corresponds to Mix, 2 corresponds to Head Mix, and 3 corresponds to Head Voice. A total of 1008 audio clips were rendered into mel-spectrograms which produced 4221 initial images before splitting and augmentation.
@@ 3.2 SVM Classification
Before experimenting with a CNN, we employed a Support Vector Machine (SVM) for audio classification, utilizing spectrograms as input. This approach was motivated by the SVM’s effectiveness in high-dimensional spaces and its ability to find optimal separating hyperplanes between classes.
We began by preprocessing the spectrogram images. The input is standardized to 154x128 pixels, applying resizing and padding as necessary. To further generalize the dataset and address potential class imbalance, we used two data augmentation techniques: horizontal flipping and brightness adjustment. Each image was flipped across the Y-axis, and two additional versions were created by adjusting the brightness to 0.8 and 1.2 times the original. We limited our approach to these techniques to balance the benefits of a larger dataset against the constraints imposed by the super-linear scaling of SVM training time.
@@
After augmentation, the spectrograms were converted to grayscale and flattened into one-dimensional vectors, resulting in a feature space of 19,712 dimensions (154x128). This transformation allowed us to represent each audio sample as a point in this high-dimensional space, where the SVM could effectively separate different vocal register classes. The combination of our targeted augmentation strategy and this high-dimensional representation provided the SVM with a rich, varied dataset that captured the characteristics of each vocal register while maintaining computational feasibility.
Figure 2: pre-processed grey-scale spectrogram example.
@@
A key advantage of our SVM approach is its ability to provide probability estimates for each class, offering insights into the model’s confidence in its predictions. This feature is particularly valuable in vocal register classification, where boundaries between classes can be subtle and gradual.
By leveraging the SVM’s strengths in handling high-dimensional data and finding optimal decision boundaries, we aimed to create a robust classifier capable of accurately distinguishing between different vocal registers based on spectrogram representations of audio samples.
@@ 3.3 CNN Architecture
The main goal of this project; however, was to employ a Convolutional Neural Network (CNN) for audio classification, leveraging its proven efficacy in image processing tasks. Like with the SVM, this approach was motivated by the transformation of audio signals into spectrograms, converting the audio classification problem into an image based task. The CNN’s ability to capture more specific patterns in spectrograms, which often correspond to specific audio features such as harmonics, or formants in vowels, made it particularly suitable for this application.
Our CNN architecture consists of three convolutional layers followed by two fully connected layers. This structure was chosen to balance depth for feature abstraction with computational efficiency. The convolutional layers, provide translation invariance – a crucial property in audio classification where important features may occur at different time points or frequency bands within the spectrogram. Following each convolutional layer, we applied max pooling to reduce spatial dimensions, introducing a degree of position invariance and managing computational complexity.
The use of ReLU activation functions after each convolutional layer introduces non-linearity, allowing the model to learn complex, non-linear decision boundaries. This is essential for capturing the intricate relationships between different audio features. The final fully connected layers combined the high-level features extracted by the convolutional layers for classification, effectively learning the optimal weighting of various audio features to discriminate between classes.
@@
By employing this CNN architecture, we aimed to leverage these advantages for effective audio classification, aiming to outperform traditional methods that rely on laborious feature engineering or simpler machine learning models. The adaptability of this architecture to audio characteristics, its ability to handle variable-length inputs (with appropriate modifications), and its potential for capturing relevant features in many audio classification tasks motivated this choice.
@@ 4.1 Training Performance
4.2 SVM Results
The SVM model demonstrated strong performance in classifying vocal registers, as shown by the metrics in Tables 1 and 2. On the training set, the model achieved near-perfect classification across all registers, with precision, recall, and F1-scores ranging from 0.98 to 1.00, and an overall accuracy of 0.99.
The confusion matrix (Table 3) provides further insight into the model’s performance. It reveals that misclassifications were most common between adjacent registers (e.g., Chest and Mix, or Mix and HeadMix), which follows with the continuous nature of vocal register transitions.
4.3 CNN Results
In the first epoch, the model started with relatively high loss values, ranging from 1.394666 to 0.632968. The validation accuracy after the first epoch was 68.0%, with an average loss of 0.787375. As training progressed, we observed a steady improvement in both training loss and validation accuracy as shown in Tables 5 and 4.
The training loss in the final epoch ranged from 0.064901 to 0.034674, showing a dramatic reduction from the initial epoch. This pattern of improvement in both training loss and validation accuracy suggests that the model effectively learned to extract relevant features from the mel-spectrograms while maintaining good generalization to unseen data. Furthermore, training was halted after the sixth epoch due to satisfactory accuracy and loss on the previous two testing sets.
@@ 5 Discussion
Overall, both models’ high performance validates our approach of using mel-spectrograms as input for vocal register classification. The success of this method suggests that mel-spectrograms effectively capture the timbral and textural differences between vocal registers, allowing the models to learn discriminative features. However, with a much more training data, and classes to identify, the CNN may prove to be more effective because it has a shown propensity for relevant feature identification. Since the SVM does not scale as well to large datasets, this CNN model may adapt better to training data containing all voice types and a wider variety of contemporary music genres.
Although the CNN slightly outperforms the SVM in overall accuracy, the SVM’s practical advantages have made it the preferred choice in the concurrently developed AVRA software interface. In this application, musicians can upload their own vocal audio, render sections of the audio into spectrograms, and then apply either the SVM or CNN for automatic analysis. Output of labels in 10 pixel increments are displayed as in figure 3
@@
Figure 3: example output of vocal audio analysis
That being said, it’s important to note some limitations of this study. The dataset was limited to contemporary male pop music, and further research would be needed to generalize these findings to other genres or voice types. This would have particularly benefited the CNN, and promoted greater functionality in the AVRA software. Additionally, while both models perform well on our dataset, real-world application would require testing on a wider range of singers and recording conditions to solidify performance. With this in mind, uture work could explore several avenues:
@@
    Investigation of model intractability, particularly for the SVM, to understand which spectral features are most important for register classification.
@@
    Comparison with other machine learning architectures, such as recurrent neural networks, which might better capture temporal dependencies in the audio signal.
@@
    Continued refinement of AVRA application for singers and vocal coaches, leveraging its practical advantages.
6 Conclusion
In this paper, we have presented two machine learning approaches to vocal register classification that show great promise, achieving high accuracy while maintaining good generalization. The SVM has shown itself to be the preferred analysis option at the present moment, but the strengths of CNNs provide a solid foundation for this task, and show promise of a more robust tool that could be leveraged across a greater variety of vocal music. Additionally, we briefly discussed the potential for real-world applications and the development of AVRA for analysis of male pop music, and how this methodology could impact vocal pedagogy and music production, providing tools for both learning and creative applications in the realm of vocal performance. As we continue to refine these models and expand their applicability, we anticipate exciting developments in the intersection of machine learning and vocal music analysis.

### v2

@@ Abstract
For singers of all experience levels, one of the most fun and daunting challenges in learning, technical repertoire is navigating placement and vocal register in and around the passagio (passage between chest voice and head voice registers). Contemporary Pop and Musical Theater solos increasingly demand strong command through and above the first passagio, and the use of various timbre and textures to achieve a desired quality. Thus, it can be difficult to identify what vocal register within the vocal range a singer is using even for advanced vocalists. This paper presents two methods for classifying vocal registers in an audio signal of male pop music through the end-to-end analysis of textural features of mel-spectrogram images. Additionally, we will discuss the practical integration of these models for vocal analysis tools, and introduce a concurrently developed software called AVRA which stands for Automatic Vocal Register Analysis. Our proposed methods achieved consistent classification of vocal register through both Support Vector Machine (SVM) and Convolutional Neural Network (CNN) models, which shows promise for robust classification possibilities across a greater range of voice types and genre.
@@ 1 Introduction
In all voices, there are two muscles responsible for manipulating the vocal folds. The thyroarytenoid (TA) muscles shorten and thicken the folds while the cricothyroid (CT) muscles lighten and lengthen. TA dominant vocal functions result in ’chest voice’ which the lower speech like vocal register in all genders. On the other hand, the CT muscles produce that pop falsetto head voice sound that has become prominent in the 21st century, and is often used to reach higher notes with a softer quality.
1.1 Interpreting Mixed Voice in Pop Music
In vocal pedagogy, ’mixed voice’ is used to denote the coordination of both TA and CT muscles, which imbues the sound with qualities from both chest voice and head voice. However, an experienced singer can change the balance of muscle function, so the term mixed voice covers the entire spectrum of chest-like to head-like timbres. For this reason, we divided the mixed voice classification into ’mix’ and ’head mix’ where mix denotes a more TA dominant, solid sound and head mix reflects more of the lighter head voice qualities.
The challenge in identifying vocal register is the variety of factors involved from person to person. In male pop music, navigating the passagio demands control over mixed voice which raises the difficulty of vocal register identification. More formally, the passagio is sequence of pitches that bridge or transition between chest and head voice. Vocal registers don’t have clear boundaries and can overlap in pitch particularly through the passagio. Besides pitch, other qualities like resonance, and placement play a role in determining the color and texture of the sound. This is how a singer can sing one note in two very different ways which is reflected in their choice of vocal register.
1.2 Advantages of Data
Driven Vocal Analysis
Expressing all these qualities of vocal audio with data was done by rendering samples of a single vocal register into mel-spectrograms. Thus, the proposed methods were achieved by analyzing the features and properties of mel-spectrograms. More specifically, the Fast-Fourier-Transform used to convert audio into spectrograms inherently increases the data from two to three dimensions. This enables for more robust automatic feature extraction via image classifiers, and improves data labeling through multi-modal identification of formants in audio clips and their corresponding images.
Figure 1: AVRA analysis output example.
The Automatic Vocal Register Analysis (AVRA) software may help students and experienced singers further their understanding and practice of contemporary vocal technique. In AVRA, the user uploads an audio clips which is then processed into a sequence of mel-spectrograms. Then user then highlights the portion of audio they want analyzed and the software returns back the spectrogram with markers and labels for vocal register shifts. The temporal visual labeling of AVRA could help vocalists see exactly where they might be struggling, and what technique their target artist is using.
@@ 2.1 Audio Classification with Images
3 Materials and
Methods
This section describes the process and pipelines used for building the dataset, pre-processing images and extracting features, and model specifications and architectures. Finally, it will conclude by summarizing the training and evaluation process that determined which models were pushed to production.
@@ 3.1 Dataset
Similarly to the paper referenced earlier, we used mel-spectrograms as the expression of the labeled audio clips. The lead vocals from WAV audio files of pop-songs were manually separated using the ’stem-splitter’ tool in Logic Pro. Then, the raw audio was imported into Audacity and rendered into mel-spectrograms with a maximum frequency of 20000 Hz, gain of 20 dB, range of 80 dB, standard Roseus color scheme, and Hann window type as the default settings. These parameters were later replicated using the Librosa library in Python for model inference pre-processing.
@@
Given that the dataset consists of 2010’s pop music, pitches ranged from C3-C5, but were generally from A3-G4. Qualitative definitions for each vocal register were agreed upon at the beginning of the project with supplementation from various vocal technique texts (Peckham 2010). Then the raw vocals of each song were split into 3 second clips and rendered into spectrograms. An image was captured for any continuously used vocal register and then later processed into a standard size for the model inputs. These sub-images were given a one-hot encoded label for one of the 4 vocal registers. In these labels, 0 corresponds to Chest Voice, 1 corresponds to Mix, 2 corresponds to Head Mix, and 3 corresponds to Head Voice. A total of 1008 audio clips were rendered into mel-spectrograms which produced 7221 initial images before splitting and augmentation.
@@ 3.2 SVM Classification
Before experimenting with a CNN, we implemented a Support Vector Machine (SVM) for audio classification, utilizing spectrograms as input. This approach was motivated by the SVM’s effectiveness in high-dimensional spaces by using the kernel trick to learn optimal separating hyperplanes between classes.
We began by pre-processing the spectrogram images. The input is standardized to 154x128 pixels, applying resizing and padding as necessary. To further generalize the dataset and address potential class imbalance, we used two data augmentation techniques: horizontal flipping and brightness adjustment. Each image was flipped across the Y-axis, and two additional versions were created by adjusting the brightness to 0.8 and 1.2 times the original. We limited our approach to these techniques to balance the benefits of a larger dataset against the constraints imposed by the super-linear scaling training time for SVM with a non-linear kernel.
@@
After augmentation, the spectrograms were converted to grayscale, which matches the pre-processing technique Lacera and Mello used. The images were then flattened into one-dimensional vectors, resulting in a feature space of 19,712 dimensions (154x128). This transformation allowed us to represent each audio sample as a point in this high-dimensional space, where the SVM could effectively separate different vocal register classes. The combination of our targeted augmentation strategy and this high-dimensional representation provided the SVM with a rich, varied dataset that captured the characteristics of each vocal register while maintaining computational feasibility.
Figure 3: pre-processed grey-scale spectrogram example of a balanced mix (label 1) spectrogram.
@@
A key advantage of our SVM approach is its ability to provide probability estimates for each class, offering insights into the model’s confidence in its predictions. This feature is particularly valuable in vocal register classification, where boundaries between classes can be subtle and gradual. By leveraging the SVM’s strengths in handling high-dimensional data we trained a classifier capable of accurately distinguishing between different vocal registers based on spectrogram representations of audio samples.
@@ 3.3 CNN Architecture
The main goal of this project; however, was to employ a Convolutional Neural Network (CNN) for audio classification, leveraging its proven efficacy in image processing tasks. Like with the SVM, this approach was motivated by the transformation of audio signals into spectrograms, converting the audio classification problem into an image based task. The CNN’s ability to capture hidden patterns in spectrograms, which often correspond to audio features such as harmonics, or formants in vowels, made it particularly suitable for this application.
Our CNN architecture consists of three convolutional layers followed by two fully connected layers. This structure was chosen to balance depth for feature abstraction with computational efficiency and generalization. The convolutional layers, provide translation invariance – a crucial property in audio classification where important features may occur at different time points or frequency bands within the spectrogram. Following each convolutional layer, we applied max pooling to reduce spatial dimensions, introducing a degree of position invariance and managing computational complexity.
The use of ReLU activation functions after each convolutional layer introduces non-linearity, allowing the model to learn, non-linear decision boundaries. This is essential for capturing the intricate relationships between different audio features. The final fully connected layers combined the high-level features extracted by the convolutional layers for classification, effectively learning the optimal weighting of various audio features to discriminate between classes.
@@
By employing this CNN architecture, we aimed to leverage these advantages for effective audio classification, aiming to outperform traditional ML methods that rely on laborious feature engineering or struggle with high dimensional data. The adaptability of this architecture to audio characteristics, its ability to handle variable-length inputs (with appropriate modifications), and its potential for capturing relevant features in many audio classification tasks motivated this choice.
@@ 4.1 Training Performance Tables
4.2 SVM Results
The SVM model demonstrated strong performance in classifying vocal registers, as shown by the metrics in Tables 1 and 2. On the training set, the model achieved near-perfect classification across all registers, with training precision, recall, and F1-scores ranging from 0.98 to 1.00, and a training of 0.99.
The confusion matrix (Table 3) provides further insight into the model’s performance. It reveals that misclassifications were most common between adjacent registers (e.g., Chest and Mix, or Mix and HeadMix). This suggests that the SVM successfully learned the features relevant to the labeled vocal registers and didn’t just memorize its training data. More importantly, this reflects the continuous nature of vocal register transitions as a fundamental structure of it’s understanding of the data.
4.3 CNN Results
@@
In the first epoch, the model started with relatively high loss values, ranging from 1.394666 to 0.632968. The validation accuracy after the first epoch was 68.0%, with a loss of 0.787375. As training progressed, we observed a steady improvement in both training loss and validation accuracy as shown in Tables 5 and 4.
The training loss in the final epoch ranged from 0.064901 to 0.034674, showing a dramatic reduction from the initial epoch. This pattern of improvement in both training loss and validation accuracy suggests that the model effectively learned to extract relevant features from the mel-spectrograms while maintaining good generalization to unseen data. Furthermore, training was halted after the sixth epoch due to prevent overfitting following strong results on the previous two testing sets.
@@ 5 Discussion
Overall, both models’ high performance validates our approach of using mel-spectrograms as input for vocal register classification. The success of this method suggests that mel-spectrograms effectively capture the timbrel and textural differences between vocal registers, allowing the models to learn discriminative features. However, with a much more training data, and classes to identify, the CNN may prove to be more effective because it has a shown propensity for relevant feature identification. Since the SVM does not scale as well to large datasets, this CNN model may adapt better to training data containing all voice types and a wider variety of contemporary music genres. With greater data training and quantity, a larger CNN has the potential to expand its domain into distinctly different types of music like classical singing, and across the full bass-soprano vocal fach system.
Although the CNN slightly outperforms the SVM in overall accuracy, the SVM’s practical advantages have made it the preferred choice in the concurrently developed AVRA software interface. In this application, musicians can upload their own vocal audio, render sections of the audio into spectrograms, and then apply either the SVM or CNN for automatic analysis. Output of labels in 10 pixel increments are marked with blue numbers (corresponding to the one-hot encoding scheme) at the bottom of the AVRA output as in figure 4.
Figure 4: example output of vocal audio analysis as shown on AVRA
@@
6 Conclusion
In this paper, we have presented two machine learning approaches to vocal register classification that show great promise, achieving high accuracy while maintaining good generalization. The SVM has shown itself to be the preferred analysis option at the present moment, but the strengths of CNNs provide a solid foundation for this task, and show promise of a more robust tool that could be leveraged across a greater variety of vocal music. Additionally, we briefly discussed the potential for real-world applications and the development of AVRA for analysis of male pop music, and how this methodology could impact vocal pedagogy and music production, providing tools for both learning and creative applications in the realm of vocal performance.
6.1 Future Considerations
Although the initial models are integrated into AVRA and performing well on this domain, it’s important to note some limitations of this study. The dataset was limited to contemporary male pop music, and further research and data would be needed to generalize these findings to other genres or voice types. This would particularly benefited the CNN, and expand the target audience of the AVRA software. Additionally, while both models perform well on our dataset, real-world application would require testing on a wider range of singers and recording conditions to solidify performance. With this in mind, future work could explore several avenues:
@@ 6.1 Future Considerations
    Investigation of the intractability of the model, particularly for the SVM, to understand which spectral features are most important for register classification.
@@
    Comparison with other machine learning architectures, such as recurrent neural network or transformer variants, which might better capture temporal dependencies in the audio signal.
@@
    Continued refinement of the AVRA application for singers and vocal coaches, leveraging its practical advantages.
As we continue to refine these models and expand their applicability, we anticipate exciting developments at the intersection of machine learning and vocal music analysis.